Astronomy Report – March 3rd

Crew 223 Astronomy Report Mar 03 2020

Name: Florian Delpech
Crew: 223
Date: 3/3/2020

MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested (choose one) MDRS-14

Objects to be Imaged this Evening: IC 342 ; NGC 2775 ; NGC 4449 ; NGC
4254 ; NGC 4125 ; NGC 3338 ; NGC 2683 ; M100 ; NGC 6503 ; NGC 5248 ; NGC
4038. This is part of the research project aiming at observing about 30
galaxies per week to try to find a supernovae.

Images submitted with this report: N/A

Problems Encountered: N/A

MUSK OBSERVATORY

Solar Features Observed: I took some images of the Sun following the
Quick guide. I did not observe any particular feature.

Images submitted with this report: A picture of the Sun

Problems Encountered: When I follow the Quick Guide, it seems that using
the different knobs should lead to see granules on the surface of the
Sun. However, I just see a quite smooth grey surface. Do I have to
adjust something?

GreenHab Report – March 3rd

Crew 223 GreenHab Report – 03-03-2020 GreenHab Officer: Valentin
BELLEMIN-LAPONNAZ

Environmental control: Ambient with Door open

Average temperature: 23.8°C
Low temperature: 19°C
High temperature: 26°C

Hours of supplemental light: 4 hours (from 9P.M to 1 a.m.)

Daily water usage for crops: 16.6 gallons

Daily water usage for research and/or other purposes: 0.25 galons

Water in Blue Tank— 201.4 gallons

Time(s) of watering for crops: 4 (2 only for tomatoes)

Changes to crops: I harvested all the Green Spinach. I’ll plant some
more plants in the pot.

Narrative: This morning, Clement PLAGNE took care of the GreenHab, as I
were on an EVA. He watered all the plants as usual. For lunch, I
harvested zucchini flowers and sage leafs to remake this delightful
Italian recipe. This afternoon, I carried out the experience Music For
Plants. I planted wheat and watercress seeds in plates. I put them in
the GreenHab and I hope they won’t dry there because of the mack of
moisture. Before dinner, I harvested the whole put of Green Spinach, as
they were starving. I also harvested some baby plants of butter crunch
lettuce to make more space in the pot.

Harvest: 3g Zucchini Flower / 3g Sage / 100g Green Spinach / 20g
ButterCrunch Lettuce

Support/supplies needed: NA

EVA Report – March 3rd

Crew 223 EVA Report 03-03-2020
EVA # 2
Author: Aurélien Mure (Commander)
Purpose of EVA: Change batteries of LOAC and MegaAres, and then first
exploration for Marion, Florian and Valentin.
Start time: 9:05 AM
End time: 11:05 AM
Narrative: First, we went out to check all the rovers and the water in
teams of two. We took then the two rovers to go on our experiments site
to change
batteries. We have had a serious problem with SpaceSuit 10 as it stoped
definitively. We reacted to this ermergency issue. We tried to restart
it and quickly we realised that we have to come back to the hab to save
me (SpaceSuit 10 user).
The EVA team worked well to make me survive! I came back to the Hab at
09:35 and I came out at 09:43 with a brand new Spacesuit. We respected
the 5min pressurisation protocol. We went back to the MegaAres location
to change the battery.
Today we needed to extract data from the instrument. We took out 2,4GB
of electric field measurement, great news for our partners from CNRS
(french national center for scientific research).
The data processing were long and after our issue with backpack 10 we
had not much time left but we went to Kissing Camel to explore during
half an hour.
We came back to the Hab at 11:00am.
Destination: Marble’s Ritual
Coordinates (use UTM NAD27 CONUS):12S 518500 4250800, 12S 518000 4249250
and 12S 518500 4249250
Participants: Aurélien Mure (EVA Leader, Commander), Valentin Bellemin
(GHO), Marion Lebrun (Scientist), Florian Delpech (Astronomer)
Road(s) and routes per MDRS Map: Entrance Road, Cow Dung Road
Mode of travel: Rovers : Spirit & Deimos

Journalist Report – March 3rd

Crew 223 Journalist Report 03Mar2020
Author: Clément Plagne, Crew Journalist
Sol 2
Title: Quick Thinking

“I’m getting no air from the suit”

“Marion, you’re taking command of this EVA”

“If this weren’t Earth I’d be dead right now”

“We need to head back to the Hab immediately”

It’s essential to have a plan – and we do. What’s even more important is
to know how to react when things start to go off the rails. Yesterday’s
EVA started off like the previous one: Aurélien, the veteran, leading
Florian, Marion, and Valentin, three rookies on their first time out.
And, much like yesterday, one of the suits began having issues with its
ventilation systems. Aurélien was the one affected, but unfortunately
this time the problem could be solved neither quickly nor easily by the
remaining crew. A routine battery replacement had suddenly turned into
an emergency situation, as the EVA leader’s helmet began fogging up from
the lack of extra air.

Outside, three Marstronauts with only minutes of EVA experience, needing
to take over from their only experienced companion. Inside, the HabCom
and the other two of us, listening in helplessly to the radio, waiting
for them to make their way back. After the cloud of confusion lifted
from the EVA team and they got control back over the situation, all
seven of us focused on how to make sure the EVA could carry on with
minimal losses. Time was indeed of the essence: there’s only so much
that can be spent outside. As two EVA members rushed in the airlock for
pressurisation, all hands were on deck to give Aurélien another suit and
help Valentin from losing his Comms headset. As quickly as they came in,
they were back out through the airlock, and the EVA carried on
nominally, only with little less time to be spent on exploration.

After a debriefing session to go over the morning’s events, time was
needed for a bit of well-deserved relaxation. Following a hearty lunch
of vegetable curry and improvised cake (powdered butter makes me sad),
we moved on to our respective experiments. Valentin was back to his
ever-relaxing GreenHab, and Florian was busy at the observatory, taking
advantage of the good weather for solar observations. Later on in the
afternoon, Blandine introduced us to a positive psychology experiment:
amongst a myriad of cards showing character strengths, we had to chose
two and explain to the others why we felt it described us well. It was a
great time for introspection, and for learning about our crewmates. Team
cohesion is paramount in missions like this one, and being reminded that
we all have great things to bring to the others is fantastic for the
rest of the mission.

Operations Report – March 3rd

Crew 223 Operations Report 03/03/2020

SOL: 3

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Monday generator:
Off at 4:30am
On at 9:10pm SOC was 83%

Tuesday generator:
Off at 5:00am
On at 9:00pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 80%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149678.53 units

Water (static tank) – 350 gallons

Static to Loft Pump used – yes

Water in GreenHab – 201.4 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes
Hours: 149.4
Beginning charge: 100%
Ending charge: 98%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: yes
Hours: 133.6
Beginning charge: 100%
Ending charge: 98%
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: no
Hours: 139.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: Run during the engineer EVA for battery
Oil Added? no
ATV Fuel Used: Nothing to report
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: The ATV 1 and 2’s oil light up when starting the ATV and
then turn off.

HabCar used and why, where? Nothing to report.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 4, 8, 9, 10, 2 were used during the
EVA. We had troubles with suit number 10 : the fans stopped after 26
minutes. We think this is because the suit was not well plugged and thus
was not charged. We are checking this and will continue tomorrow.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Following Scott Davis’ recommendations, we
tried to fix suit 7. We checked most of the electrical connection points
and it doesn’t seem to be a loose electrical connection. We tend for a
problem with the suit’s battery (but we need to confirm this guess
tomorrow), and we will also remove the fans to make sure they are not
obstructed.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Do you know were is
the key of the upper black storage box in the RAM (I would like to open
the drawer containing the wire strippers) ? I only found the one opening
the lower black storage box.